Write an equation of the hyperbola with vertices at (–7, 0) and (7, 0) and foci at (–9, 0) and (9, 0)
Answer by lwsshak3(6463)

Write an equation of the hyperbola with vertices at (–7, 0) and (7, 0) and foci at (–9, 0) and (9, 0)
This is a hyperbola with horizontal transverse axis.
Its standard form of equation: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center.
For given hyperbola:
center: (0,0)
length of horizontal transverse axis=14=2a
a=7
a^2=49
..
foci
c=9
c^2=81
..
c^2=a^2+b^2
b^2=c^2-a^2=81-49=32
b=√32≈5.6
..
equation:
x^2/49-y^2/32=1